OpenStudy (anonymous):

Compute the cube root of -729.

OpenStudy (anonymous):

is that negative 729?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that's not true it's -9

OpenStudy (anonymous):

you can take odd roots of - numbers

OpenStudy (anonymous):

Read square root, my bad.

OpenStudy (anonymous):

One of my selections was the cube of -729 does not exist. So that's it thanks.

OpenStudy (anonymous):

no it does exist

OpenStudy (anonymous):

Prime factors of 729 are 3^6. The cube root of -729 therefore is -(3^6)^1/3 =-3^2 =-9